40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Dec 7 13:11:57 2014, manasa, Summary, General, Frequency Offset Locking - To Do List FOLtodolist.pdf
    Reply  Mon Dec 8 15:54:39 2014, manasa, Summary, General, Dec 8 - Check Frequency Counter module 
       Reply  Mon Dec 8 20:53:51 2014, diego, Summary, General, Dec 8 - Check Frequency Counter module 
          Reply  Tue Dec 9 16:06:46 2014, diego, Summary, General, Dec 8 - Check Frequency Counter module 
    Reply  Tue Dec 9 00:30:27 2014, manasa, Summary, General, Dec 9 - Elaborate to do list FOL_FiberChassis.pdf
       Reply  Tue Dec 9 16:07:16 2014, manasa, Summary, General, Dec 9 - FC module and fiber chassis FCmodule.png
    Reply  Wed Dec 10 16:12:29 2014, manasa, Summary, General, Dec 10 - PSL table FiberMod_front.jpgFOL_fiber.jpg
       Reply  Thu Dec 11 12:50:12 2014, manasa, Summary, General, PSL table optical layout 
          Reply  Thu Dec 11 23:34:06 2014, manasa, Summary, General, PSL table optical layout 
             Reply  Fri Dec 12 02:30:25 2014, Jenne, Summary, General, PSL table optical layout 
    Reply  Fri Dec 12 14:38:39 2014, manasa, Summary, General, Frequency Offset Locking - To Do List (Revised) FOLtodolist.pdf
       Reply  Fri Dec 12 15:19:04 2014, manasa, Summary, General, Dec 12 - PSL table PSLtoFiber.pngPSLfiberChassis.pngPSLfiberChassis.png
       Reply  Wed Dec 17 11:57:56 2014, manasa, Summary, General, Y arm optical layout 58.png
Message ID: 10792     Entry time: Fri Dec 12 15:19:04 2014     In reply to: 10791
Author: manasa 
Type: Summary 
Category: General 
Subject: Dec 12 - PSL table 

Quote:

Unfortunately the order placed for beam samplers last week did not go through. These will be used at the X and Y end tables to dump the unwanted light appropriately. Since they will not be here until Tuesday, I revised the timeline for FOL related activities accordingly.

I was working on the PSL table today. 

Since the rejected 1064nm light after the SHG crystal is not easily reachable to measure beam widths close to the waist, I put a lens f=300mm and measured the beam size around its focus. I used this data and redesigned the telescope using 'a la mode'.

I used a beam splitter to attenuate the beam directed towards the fiber. The reflected beam from BS has been dumped (I need to find a better beam dump than what is being used right now. 

I have only ~200uW at the input of the fiber coupler after the BS and 86uW at the output of the fiber (43% coupling)

I moved the GTRY DC photodiode and the lens in front of it to make space for the fiber coupler mount.

The layout on the PSL table right now is as shown below.

PSLtoFiber.png 

I have also put the fiber chassis inside the PSL enclosure on the rack. I moved the coherent spectrum analyser controller that is not being used to make space on the rack.

PSLfiberChassis.png

 
Attachment 2: PSLfiberChassis.png  1.256 MB  | Hide | Hide all
PSLfiberChassis.png
ELOG V3.1.3-